how much energy must be absorbed by a 150 g sample of ice at 0.0 c that melts and then warms to 25.0c
Answer #1

The heat absorbed by water = m*s*DT+n*DHfus

                                               = 150*4.184*(25-0) + (150*334) = 65790 joules

The heat energy must be absorbed by a 150 g sample of ice at 0.0 c that melts and then warms to 25.0c is 65790 joules or 65.790 KJ